# Fast In-Spectrum Graph Watermarks

### v1

@@ Abstract
We address the problem of watermarking graph objects, which consists in hiding information within them, to prove their origin. The two existing methods to watermark graphs use subgraph matching or graph isomorphism techniques, which are known to be intractable for large graphs.
To reduce the operational complexity, we propose FFG, a new graph watermarking scheme adapted from an image watermarking scheme, since graphs and images can be represented as matrices. We analyze and compare FFG, whose novelty lies in embedding the watermark in the Fourier transform of the adjacency matrix of a graph. Our technique enjoys a much lower complexity than that of related works (i.e. in $\mathcal{O}\left(N^{2}\log N\right)$), while performing better or at least as well as the two state-of-the-art methods.
@@ I Introduction
Creating graphs requires extensive data collection and pre-processing, making them highly valuable assets. Graphs emerge from a wide range of applications, including webs of trust in blockchains, circles in social networks, and co-purchasing networks in e-commerce platforms. The creators of these graphs often seek recognition, similar to the practices encouraged by Creative Commons licences. Watermarking schemes can satisfy this demand by embedding provenance information as a digital signature that can be extracted later to prove ownership.
Despite its potential, the watermarking of graph objects remains uncommon. None of the popular online platforms that share graph representations [24, 22, 32] currently propose watermarked graphs. This is mainly due to the limited number of existing techniques that are too computationally expensive. These techniques rely on intensive graph-related operations, such as subgraph matching and isomorphism [13, 38].
In this paper, we introduce a competitive watermarking scheme inspired by image watermarking techniques [4]. By treating adjacency matrices of graphs analogously to pixel matrices in images, we propose to leverage complexity-reduced computational operations for watermarking graph adjacency matrices. An overview of methods for watermarking graph-related objects is provided in the appendix B.
Contributions. Our contributions are as follows.
1) We introduce a novel analytical framework that applies an image processing scheme directly to the adjacency matrices of graphs, with the aim of using operations faster than the NP-complete related complexities of the related works [13, 38].
2) The main technical challenge in applying an image watermarking technique in a graph context is related to the binary nature of an edge’s presence in an adjacency matrix, as opposed to real numbers in an image matrix. This forces binarization. Our formal analysis studies the necessary binarization of the watermark to verify that it effectively leads to a unique watermark in the graph (i.e. without collision with another watermark insertion), setting the stage for the subsequent practical applications.
3) Finally, we perform a head-to-head comparison of FFG with the related work in the graph domain (schemes by Zhao et al. [38] and Eppstein et al. [13]). We conclude that FFG performs at least as well as its competitors, with the advantage of being much faster.
@@ II Goals and Threat Model
The following section describes watermarking schemes and the goals of such methods.
@@ II-A Watermarking Graphs: Threat Model
A graph watermarking scheme is a set of functions (Keygen, Embed, Extract), applied to a graph [13]. With Keygen, the owner of a graph generates a secret key. This key is embedded into her graph $G$ with the Embed operation. It produces a modified graph $G_{W}$ and a watermark $W$ (which is related to the difference between the original and the watermarked graph). The graph’s owner can prove her ownership of a suspected shared graph $G^{*}$ with the watermark $W$ using the Extract function. Extract returns a success if $G^{*}$ is watermarked with $W$. Otherwise, either 1) $G^{*}$ was not watermarked with $W$ and the considered watermarking scheme (or not watermarked at all), or 2) an attacker who wanted to prevent extraction has modified $G^{*}$ enough to prevent Extract from succeeding.
@@ II-B Design Goals for Watermarking Schemes
The Embed function must be injective. In other words, a graph watermarked independently with two different keys should not produce two identical graphs.
@@
A robust watermarking scheme must be resistant to modifications (i.e. attacks) performed on the watermarked graph. More specifically, a key extraction (Extract) must work with a high probability, despite attacks.
Other goals such as undetectability and distortion effect are defined and studied in Appendix E. These are additional goals we set with respect to the state-of-the-art watermarking methods.
III FFG: Cox et al. Scheme for Graphs
This work considers undirected and unweighted graphs, in order to compete with state-of-the-art schemes [13, 38] described in Section V-A. Since the goal is to exploit watermarking schemes originally designed for images, we assume that we can label the input graph arbitrarily, so that a deterministic mapping of its vertices in the adjacency matrix can be performed [14, 17]. This assumption is realistic in many cases where vertices are naturally labeled and it is consistent with the fact that for images, pixels have a deterministic position in their matrix representation.
@@ III-A Framework Rationale
Notations introduced hereafter are also summed up in Table III and Figure 7 (Appendix A).
@@
Therefore, our framework includes additional steps to transform $A^{\prime}$ back into an undirected and unweighted graph. Part of our challenge is to analyze the consequences of these steps. More specifically, the required binarization is the thresholding of all values in $A^{\prime}$ by the average of $A$. This real value is denoted $av(A)$. Each element is set to 00 if its modulus is less than the average of the original adjacency matrix $A$, otherwise $1$.
@@
Next, we describe the original scheme of Cox et al. for watermarking images, before detailing its adaptation to graphs in our framework.
@@
In [4], the watermark is a sequence of a fixed number of random reals chosen from a Gaussian distribution. The authors compute the Fourier transform of the original image and insert the watermark into the largest low-frequency coefficients of the transformed matrix. They propose three ways to insert the watermark into the highest magnitude coefficients, including a simple additive method.
For a given image (possibly modified), a watermark is extracted in the spectral domain and compared with the original watermark. A similarity score is computed, which decides whether this image is watermarked or not.
III-C The FFG Scheme for Watermarking Graphs
We now propose FFG as an adaptation of the Cox et al. scheme to be applicable to unweighted graphs. We first present the three functions Keygen, Embed and Extract. The pseudocode is given in Algorithm 1. We then discuss the main differences with the original Cox et al. scheme.
Algorithm 1 FFG scheme
1:function Keygen($m,\sigma$)
@@ III-C The FFG Scheme for Watermarking Graphs
Keygen$(m,\sigma)$ generates a key as a Gaussian vector of $m$ values with standard deviation $\sigma$. This step is identical to that of Cox et al..
@@
The graph owner wants to embed a key $\omega$ generated with the function Keygen into her graph $G$ thanks to its adjacency matrix $A$. First, the sequence $\chi=\{(i_{1},j_{1}),\dots,(i_{m},j_{m})\}$ is computed as the sequence of the first $m$ indices that would sort the modulus of the Fourier coefficients of $A$. The sequence $\chi$ is used to compute the intermediate watermark $\mathcal{W}$ of the graph owner in the Fourier spectrum. $\mathcal{W}$ is the matrix representation of the key where each element $\omega[n]$ is at position $\chi[n]=(i_{n},j_{n})$. Then, the Fourier inverse of the intermediate watermark $\mathcal{W}$ is added to the adjacency matrix $A$ to obtain a real matrix $A^{\prime}$. The algorithm continues with binarization and symmetrization to obtain $A_{W}$. The watermark $W$ is defined as the difference of all coefficients of the Fourier transform of $A$ and $A_{W}$: $W=FT(A)-FT(A_{W}).$
@@
The suspected watermark $W^{*}$ is compared to the original watermark $W$ using the $2$-norm of their difference. If the difference between $W^{*}$ and $W$ is less than a threshold, the extraction is hypothetically successful.
The specificity of FFG
The first difference between the original scheme of Cox et al. [4] and FFG is that the watermark key cannot be chosen arbitrarily, but instead depends on the binarization operation. What is left after binarization depends on the chosen length $m$ and the standard deviation $\sigma$ of the elements in that key. Recommendations in Cox et al. [4] are no longer valid due to this binarization step. Section IV-C discusses how to properly obtain a key.
The second specificity is also due to the binarization step. This step causes the key to be spread over the entire transformed matrix. As a consequence: $\mathcal{W}\neq W.$ Thus, the watermark is defined over all matrix coefficients and not only over the ones specified by $\chi$. The function Extract is applied to the watermark $W$ resulting from the binarization after embedding $\omega$ with the Cox et al. scheme, and not applied directly to the key $\omega$ in the coefficients specified by $\chi$. Given the key $\omega$, the watermark $W$ can be derived by the embedding function or by calling Extract before comparing to $W^{*}$.
@@ III-D Scheme complexity vs related works
For comparison, the scheme from Zhao et al. leverages the NP-complete subgraph matching routine. They argue that the computational complexity is lower on graphs with a ”very high level of node heterogeneity” in terms of degrees, without providing a complexity analysis for this claim (see [38], Section 4); the problem still remains NP-complete in the general case. While Eppstein et al. [13] prove the correctness of their scheme for Erdős-Rényi and power-law graph families, they do not provide a bounded complexity analysis for these, and thus also remain in the NP-complete complexity class due to the subgraph isomorphism routine. All three schemes are compared in terms of their practical execution times in Section V-B.
@@ IV Experiments: Compliance to Design Goals
This section presents the results of an extensive experimental study of FFG. It is organized to shed light on the practical compliance of FFG to the design goals detailed in Section II-B.
@@ IV-A Experimental Setup
Experiments are performed using Python 3. Reported timings are obtained from execution on a server equipped with Intel Xeon E5-2630 v3 CPUs, running at 2.40GHz.
@@
For experiments that require the variation of a graph property, graph generators from the Python package NetworkX [16] are used. These are well-known graph models, that have been used to model various real-world graphs: Erdős-Rényi (hereafter referred to as ER), Barabási-Albert (BA), or Watts-Stroggats (WS) models.
@@
Experiments with FFG are also performed on eight real-world large graphs with different edge densities, obtained from the SNAP repository [24] (e.g. Pokec) and the NetworkRepository [32] (e.g. Flickr). These graphs have different structures, which makes them interesting to study.
@@
The properties and information to preserve are highly context-dependent, depending on the intended use of a graph. It is impractical to cover all possible scenarios. Due to its generality and widespread use in the state-of-the-art [13, 38], we consider the graph edit distance, as a metric to quantify the degradation of a graph after these attacks:
Definition.
The graph edit distance (denoted ED) is the percentage of edges that distinguish two graphs. In the sequel, the distance is measured relative to the original graph.
For instance, if the $embed$ function flips $200$ edges from a graph $G$ containing $2,000$ edges, then the distance based on the fraction of edges between $G^{\prime}$ and $G$ is $10\%$ edges. We emphasize that the ED can be greater than $100$% because we can add edges and not just remove them. In fact, one can flip at most $N(N-1)/2$ edges from the input graph.
@@ IV-B Adaptations from the Theoretical Scheme
IV-B1 FFG Scalability
To increase the scalability of their schemes when applied to large graphs, both related works [13, 38] filter out the input graph to operate on a smaller set of vertices. Inspired by the Cox et al. scheme [4] FFG also performs a dimensionality reduction. This step is described in detail in the Appendix D.
We empirically evaluate the scalability of FFG by increasing the size of the input graph from $N=10^{3}$ to $10^{7}$. To maintain a similar graph structure as the size increases, a BA graph generator is used. Its attachment parameter is set to $3$ (i.e. each new vertex attaches to $3$ existing vertices according to preferential attachment). The length of the key is arbitrarily always the same ($m=200$) because all the graphs have the same attachment parameter (see Section IV-C).
The measured timings are as follows. For embedding: $0.20,~{}18,~{}20,$ $46$ and $358$ seconds for BA with sizes $1k,~{}10k,~{}100k,~{}1M$ and $10M$ respectively. For extraction, and the same graph sizes: $0.13,~{}15,~{}16,~{}17$ and $20$ seconds. The execution time of the key generation is almost constant at $0.8$ milliseconds. These observations are fully consistent with the timing complexity analysis, and therefore confirm that our FFG is tractable for large graphs.
@@ IV-B2 Resilience to Attacks and the Similarity Threshold
The ideal case: the watermarked graph is never attacked.
We demonstrate that if attacks on $G_{W}$ are not considered, $\theta=0$ can be set with the guarantee that the watermark of the graph’s owner will always be extracted. That is, considering unattacked watermarked graphs, no false negatives can occur.
Proof. A false negative occurs if Extract returns a failure even though the tested graph was indeed watermarked with a given watermark. In particular, the extraction will fail only if the following condition is not satisfied: $\left\|W^{*}-W\right\|_{2}\leq\theta*\left\|W\right\|_{2},$ where $W^{*}$ and $W$ are computed according to Algorithm 1, and while the tested graph $G^{*}$ is the watermarked graph $G_{W}$, that has not been attacked. Thanks to the deterministic mapping of vertices to adjacency matrices, $A^{*}=A_{W}$. Since the Fourier transform is linear, the condition $\left\|W^{*}-W\right\|_{2}\leq\theta*\left\|W\right\|_{2},$ becomes: $0\leq\theta*\left\|W\right\|_{2}$, which always holds since $\theta\geq 0$. Thus if $\theta=0$, no false negatives can occur with the FFG scheme.
A realistic case where the watermarked graph is attacked
In a real scenario, after the owner of a graph shares her watermarked graph, it can be modified. A false negative occurs when Extract returns a failure, despite the tested graph was indeed watermarked with the owner’s watermark. In this case, $\theta>0$ defines the similarity above which the graph’s owner considers the extracted watermark to be close enough to the original one, despite modifications.
We set $\theta$ to tolerate $10$% edge flips, which is clearly a conservatively high value (according to related work [13, 38], as and it will be illustrated in Section IV-C). Consequently, no false negatives can occur with less than $10$% edge flips. The effect of $\theta$ on false positives is studied in Section IV-E.
@@
We now experiment with the distortion caused by embedding the watermark in a given graph. We study it theoretically in Appendix C.
Practical guidelines for low distortion
From Section II-B, low distortion is targeted for the watermark embedding. According to related work [13, 38], the ED of large graphs must be at most $10^{-2}$edges. We now give a method to automatically set the parameters suitable for watermarking with such a minimal distortion.
Figure 1 illustrates the link between $\sigma$ and $m$ for our three generative graph models, and a small distortion of $0.005\%$ edges.
Key lengths used in [13] depend on densities (being defined as the number of edges over the number of vertices) [26], while those of [38] depend on the number of vertices. We set the key length $m$ of FFG as the key length used in [13] because the experiment in Section IV-D shows that FFG is also related to densities.
Once the key length $m$ is set as in [13], it is always possible to find a value of $\sigma$ to watermark the graph while achieving the distortion goal. Experiments have shown that for any graph $G$ and for any fixed key length $m\in[\![1,N_{0}*(N_{0}-1)/2]\!]$, there exists a value $\sigma_{max}$ to watermark the graph with a strictly positive edit distance. An automated way to set $\sigma$ under the small ED constraint is to use a dichotomous search in $[\![1,\sigma_{max}]\!]$. As shown in Appendix D, the proposed guidelines do not affect the scalability of FFG.
@@
The three large and real graphs are watermarked using this method, leading to the parameters in Table I. Note that the values of $m$ differ significantly between graphs, depending on their inner topological properties. The resulting EDs are effectively kept well below the target of $10^{-2}$%. Goal 1 is thus achieved.
Further experiments on the impact of FFG on graphs are studied in Appendix E.
@@
To assess the uniqueness property resulting from FFG, two keys $\omega_{1}$ and $\omega_{2}$ are generated with the same parameters ($m$ and $\sigma$) found by the automatic dichotomous procedure. There is a lack of uniqueness if the two keys result in the same watermark (i.e. if $\omega_{1}\neq\omega_{2}$ and $G_{W_{1}}=G_{W_{2}}$).
We experiment with uniqueness for the three graph models and various graph densities in Figure 2. There are some collisions for the lowest densities, beyond $2$. This is not an issue per se, as these graphs are rare or atypical: a density less than $2$ means that a graph has, or is close to having, disconnected components. Above this threshold, there is no collision; this is the expected setup, since collected graphs generally have high densities (e.g., from 12 to 36 in web crawls [26], or without density limit for graphs studied in node classification, where edges are even removed to reduce the density before processing [25]). Goal 2 is thus validated for practical setups.
@@
We now study the effect of $\theta$ on false positives.
Let $G_{W}$ be the watermarked graph of $G$ with the key $\omega$, and $G^{*}$ be any other graph not watermarked with $\omega$. Recall that Algorithm 1 with $G^{*}$ results in a false positive if the extraction of $W$ in $G^{*}$ succeeds.
@@
The goal of the following experiment is to determine the value of $\theta$ for which false positives occur when attempting an extraction on non-watermarked graphs. Ten BA graphs are generated with density $\frac{|E|}{|V|}=5$ (according to the uniqueness experiment in Figure 2). For each of them, we first inserted a key $\omega$; then we re-generated five more graphs with the same generative parameters, and tried to extract $\omega$ from them. Extractions are performed for different values of $\theta$. Figure 3 shows that as long as $\theta$ is less than or equal to $5$, there are no false positives. Then, values of $\theta$ above $6$ introduce some. Thus, $\theta=1.4$ is consistent with the $2$-norm of the watermark to tolerate $10$% edge flips for a graph of density $5$ in BA models. Goal 3 is then achievable by parameterization.
In the remaining experiments, $\theta$ is set to tolerate $10$% edge flips since no false positives or false negatives occur in this setup.

### v2

# Fast&Fourier: In-Spectrum Graph Watermarks

@@ Abstract
Graphs commonly represent complex interactions, and protecting their provenance helps prevent misuse and preserves trust in the data’s source. We address the problem of watermarking graph objects, which involves hiding information within them to prove their origin. The two existing methods for watermarking graphs employ subgraph matching or graph isomorphism techniques, which are known to be intractable for large graphs. To reduce operational complexity, we design a new graph watermarking scheme, F&F, based on an image watermarking approach that leverages the fact that graphs and images share matrix representations. We analyze and compare F&F, whose novelty lies in embedding the watermark in the Fourier transform of the adjacency matrix of a graph. Our technique enjoys a significantly lower complexity than that of related works (i.e., in $\mathcal{O}\left(N^{2}\log N\right)$), while performing at least as well as, if not better than, the two state-of-the-art methods.
@@ I Introduction
Creating graphs requires extensive data collection and pre-processing, making them highly valuable assets. Graphs emerge from a wide range of applications, including webs of trust in blockchains, circles in social networks, and co-purchasing networks in e-commerce platforms. The creators of these graphs often seek recognition, similar to the practices encouraged by Creative Commons licences. Watermarking schemes meet this demand by embedding provenance information as a digital signature, which the graph owner can extract later to prove ownership.
Despite its potential, the watermarking of graph objects remains uncommon. None of the popular online platforms that share graph representations [24, 22, 32] currently propose watermarked graphs. The main reason is the limited number of existing techniques, which are too computationally expensive. These techniques rely on intensive graph-related operations, such as subgraph matching and isomorphism [13, 38].
In this paper, we introduce a competitive watermarking scheme inspired by image watermarking techniques [4]. By treating adjacency matrices of graphs analogously to pixel matrices in images, we propose to leverage complexity-reduced computational operations for watermarking graph adjacency matrices. The appendix -B presents an overview of methods for watermarking graph-related objects. Our contributions are:
1) We introduce a novel analytical framework that applies an image processing scheme directly to the adjacency matrices of graphs, to use operations faster than the NP-complete related complexities of the related works [13, 38].
2) The main technical challenge in applying an image watermarking technique in a graph context is related to the binary nature of an edge’s presence in an adjacency matrix, as opposed to real numbers in an image matrix. This forces binarization. Our formal analysis examines the necessary binarization of the watermark to verify that it effectively results in a unique watermark in the graph (i.e., without collision with another watermark insertion), thereby setting the stage for subsequent practical applications.
3) Finally, we perform a head-to-head comparison of F&F with the related work in the graph domain (schemes by Zhao et al. [38] and Eppstein et al. [13]). We conclude that F&F performs at least as well as its competitors, with the advantage of being much faster.
@@ II-A Watermarking Graphs: Threat Model
A graph watermarking scheme is a set of functions (Keygen, Embed, Extract), applied to a graph [13]. With Keygen, the owner of a graph generates a secret key. This key is embedded into her graph $G$ with the Embed operation. It produces a modified graph $G_{W}$ and a watermark $W$ (which is related to the difference between the original and the watermarked graph). The graph’s owner can prove her ownership of a suspected shared graph $G^{*}$ with the watermark $W$ using the Extract function. Extract returns a success if $G^{*}$ is watermarked with $W$. Otherwise, either 1) $G^{*}$ does not contain the watermark $W$ embedded using the considered scheme (or does not contain any watermark), or 2) an attacker has modified $G^{*}$ sufficiently to prevent Extract from succeeding.
@@ II-B Design Goals for Watermarking Schemes
The Embed function must be injective: a graph watermarked independently with two different keys should not produce two identical graphs.
@@
A watermarking scheme must be resistant to modifications (i.e., attacks) performed on the watermarked graph. More specifically, the key extraction must operate with a high probability, even in the presence of attacks.
Other goals, such as undetectability and distortion effect, are defined and studied in Appendix -E. We have established a framework based on the state-of-the-art watermarking methods.
III F&F: Cox et al. Scheme for Graphs
This work considers undirected and unweighted graphs, to compete with state-of-the-art schemes [13, 38] described in Section V-A. Since the goal is to exploit watermarking schemes designed initially for images, we assume that we can label the input graph arbitrarily, so that a deterministic mapping of its vertices in the adjacency matrix can be performed [14, 17]. This assumption is realistic in many cases where vertices are naturally labeled, and it is consistent with the fact that for images, pixels have a deterministic position in their matrix representation.
@@ III-A Framework Rationale
Therefore, our framework includes additional steps to transform $A^{\prime}$ back into an undirected and unweighted graph. A key part of our challenge is to analyze the consequences of these steps. More specifically, the required binarization involves thresholding all values in $A^{\prime}$ by the average of $A$. Let $av(A)$ denote this real value. Each element becomes 0 if its modulus is less than the average of the original adjacency matrix $A$; otherwise, it becomes 111.
@@
Next, we describe the original scheme of Cox et al. for watermarking images.
@@
In [4], the watermark is a sequence of a fixed number of random reals chosen from a Gaussian distribution. The authors compute the Fourier transform of the original image and insert the watermark into the most significant low-frequency coefficients of the transformed matrix. They propose three ways to insert the watermark into the highest magnitude coefficients, including a simple additive method.
Starting from a possibly altered image, the procedure extracts a watermark in the spectral domain, compares it to the original, and computes a similarity score to decide if the image contains the watermark.
III-C The F&F Scheme for Watermarking Graphs
We now propose F&F as an adaptation of the Cox et al. scheme for unweighted graphs with pseudo-code Algorithm 1. We discuss the main differences with the original Cox et al. scheme.
Algorithm 1 F&F scheme
[…]
@@ III-C The F&F Scheme for Watermarking Graphs
[…]
@@
The graph owner wants to embed a key $\omega$ generated with the function Keygen into her graph $G$ thanks to its adjacency matrix $A$. First, the sequence $\chi=\{(i_{1},j_{1}),\dots,(i_{m},j_{m})\}$ is computed as the sequence of the first $m$ indices that would sort the modulus of the Fourier coefficients of $A$. The process relies on the sequence $\chi$ to compute the intermediate watermark $\mathcal{W}$ of the graph owner in the Fourier spectrum. $\mathcal{W}$ is the matrix representation of the key where each element $\omega[n]$ is at position $\chi[n]=(i_{n},j_{n})$. Then, the Fourier inverse of the intermediate watermark $\mathcal{W}$ is added to the adjacency matrix $A$ to obtain a real matrix $A^{\prime}$. The algorithm continues with binarization and symmetrization to obtain $A_{W}$. The watermark $W$ is equal to the difference of all coefficients of the Fourier transform of $A$ and $A_{W}$: $W=FT(A)-FT(A_{W}).$
@@
The suspected watermark $W^{*}$ is compared to the original watermark $W$ using the 222-norm of their difference. If the difference between $W^{*}$ and $W$ is less than a threshold, the extraction is hypothetically successful.
The specificity of F&F
The first difference between the original scheme of Cox et al. [4] and F&F concerns the watermark key, which depends on the binarization operation rather than allowing arbitrary selection. What is left after binarization depends on the chosen length $m$ and the standard deviation $\sigma$ of the elements in that key. Recommendations in Cox et al. [4] are no longer valid due to this binarization step. Section IV-C discusses how to obtain a key properly.
The second specificity is also due to the binarization step. This step spreads the key across the entire transformed matrix. As a consequence: $\mathcal{W}\neq W.$ Thus, the watermark is defined over all matrix coefficients and not only over the ones specified by $\chi$. The function Extract is applied to the watermark $W$ resulting from the binarization after embedding $\omega$ with the Cox et al. scheme, and not applied directly to the key $\omega$ in the coefficients specified by $\chi$. The procedure uses the key $\omega$ to derive $W$, either by computing it through the embedding function or by calling Extract, then comparing it to $W^{*}$.
@@ III-D Scheme complexity vs related works
For comparison, Zhao et al. use an NP-complete subgraph matching routine and suggest lower complexity on graphs with ”very high node degree heterogeneity”, though without formal analysis (see [38], Section 4); the problem remains NP-complete in general. Eppstein et al. [13] prove correctness for Erdős-Rényi and power-law graphs but provide no bounded complexity analysis, so their scheme also remains NP-complete. Section V-B compares all three schemes by execution time.
@@ IV Experiments: Compliance to Design Goals
This section presents the results of an extensive experimental study of F&F. The structure highlights how F&F meets the design goals outlined in Section II-B.
@@ IV-A Experimental Setup
Using Python 3, we conduct experiments and record timings from runs on a server featuring Intel Xeon E5-2630 v3 CPUs operating at 2.40 GHz.
@@
For experiments involving variations in graph properties, we use graph generators from the Python package NetworkX [16]. These generators implement well-known models applied to various real-world graphs, such as the Erdős-Rényi (ER), Barabási-Albert (BA), and Watts-Strogatz (WS) models.
@@
We also conduct experiments with F&F on eight real-world large graphs with varying edge densities and different structures, obtained from the SNAP repository [24] (e.g., Pokec) and the Network Repository [32] (e.g., Flickr).
@@
The properties and information to preserve are highly context-dependent, depending on the intended use of a graph. It is impractical to cover all possible scenarios. Due to its generality and widespread use in the state-of-the-art [13, 38], we consider the graph edit distance to quantify the degradation of a graph:
Definition
The graph edit distance (denoted ED) is the percentage of edges that distinguish two graphs. In the sequel, the distance refers to the comparison with the original graph.
For instance, if the $embed$ function flips 200200200 edges from a graph $G$ containing 2,0002,0002 , 000 edges, then the distance based on the fraction of edges between $G^{\prime}$ and $G$ is 101010 % edges. We emphasize that the ED can be greater than 100100100 % because we can add edges, not just remove them. One can flip at most $N(N-1)/2$ edges from the input graph.
@@ IV-B Adaptations from the Theoretical Scheme
IV-B1 F&F Scalability
To increase the scalability of their schemes when applied to large graphs, both related works [13, 38] filter out the input graph to operate on a smaller set of vertices. Inspired by the Cox et al. scheme [4] F&F also performs a dimensionality reduction. This step is described in detail in the Appendix -D.
@@ IV-B2 Resilience to Attacks and the Similarity Threshold
The ideal case: no attack on the watermarked graph
We show that when attacks on $G_{W}$ do not occur, setting $\theta=0$ guarantees the successful extraction of the graph owner’s watermark. In other words, no false negatives arise for unattacked watermarked graphs.
Proof. A false negative occurs when Extract fails despite the tested graph containing the given watermark. Specifically, the extraction fails only if the following condition does not hold: $\left\|W^{*}-W\right\|_{2}\leq\theta*\left\|W\right\|_{2},$ where $W^{*}$ and $W$ are computed according to Algorithm 1, and while the tested graph $G^{*}$ is the watermarked graph $G_{W}$, that has not been attacked. Thanks to the deterministic mapping of vertices to adjacency matrices, $A^{*}=A_{W}$. Since the Fourier transform is linear, the condition $\left\|W^{*}-W\right\|_{2}\leq\theta*\left\|W\right\|_{2},$ becomes: $0\leq\theta*\left\|W\right\|_{2}$, which always holds since $\theta\geq 0$. Thus, if $\theta=0$, no false negatives can occur with the F&F scheme.
A realistic case with an attacked watermarked graph
In practical scenarios, the owner shares a watermarked graph that may undergo modifications. A false negative occurs when Extract fails despite the graph containing the owner’s watermark. Here, $\theta>0$ sets the similarity threshold above which the owner accepts the extracted watermark as sufficiently close to the original, despite alterations.
We set $\theta$ to tolerate 101010% edge flips, which is a conservatively high value (according to related work [13, 38], as it will be illustrated in Section IV-C). Consequently, no false negatives can occur with less than 101010 % edge flips. We study the effect of $\theta$ on false positives in Section IV-E.
@@
According to related work [13, 38], the ED of large graphs must be at most $10^{-2}$edges. We now present a method for automatically setting parameters suitable for watermarking with minimal distortion. We study it theoretically in Appendix -C.
Figure 1 illustrates the link between $\sigma$ and $m$ for our three generative graph models, and a slight distortion of 0.0050.0050.005% edges.
Key lengths used in [13] depend on densities (being defined as the number of edges over the number of vertices) [26], while those of [38] depend on the number of vertices. We set the key length $m$ of F&F as the key length used in [13] because the experiment in Section IV-D shows that F&F is also related to densities.
Given the key length $m$ defined as in [13], it is always possible to find a value of $\sigma$ to watermark the graph while achieving the distortion goal. Experiments have shown that for any graph $G$ and any fixed key length $m\in[\![1,N_{0}*(N_{0}-1)/2]\!]$, there exists a value $\sigma_{max}$ to watermark the graph with a strictly positive edit distance. An automated way to set $\sigma$ under the small ED constraint is to use a dichotomous search in $[\![1,\sigma_{max}]\!]$. As shown in Appendix -D, the proposed guidelines do not affect the scalability of F&F.
@@
Applying this method to the three significant real graphs results in the parameters listed in Table I. The key length mmm varies significantly according to each graph’s inner topological characteristics. The observed edit distances remain well under the $10^{-2}$target, fulfilling Goal 1.
Further experiments on the impact of F&F on graphs appear in Appendix -E.
@@
The uniqueness property resulting from F&F is assessed by generating two keys $\omega_{1}$ and $\omega_{2}$ with the same parameters ($m$ and $\sigma$) found by the automatic dichotomous procedure. There is a lack of uniqueness if the two keys result in the same watermark (i.e., if $\omega_{1}\neq\omega_{2}$ and $G_{W_{1}}=G_{W_{2}}$).
We experiment with uniqueness for the three graph models and various graph densities in Figure 2. There are some collisions for the lowest densities, beyond 222. It is not an issue per se, as these graphs are rare or atypical: a density less than 222 means that a graph has, or is close to having, disconnected components. Above this threshold, there is no collision; this is the expected setup, since collected graphs generally have high densities (e.g., from 12 to 36 in web crawls [26], or without density limit for graphs studied in node classification, where edges are even removed to reduce the density before processing [25]). It thus fulfills Goal 2 for practical setups.
@@
We now study the effect of $\theta$ on false positives. Let $G_{W}$ be the watermarked graph of $G$ with the key $\omega$, and $G^{*}$ be any other graph not watermarked with $\omega$. Recall that Algorithm 1 with $G^{*}$ results in a false positive if the extraction of $W$ in $G^{*}$ succeeds.
@@
The goal of the following experiment is to determine the value of $\theta$ for which false positives occur when attempting an extraction on non-watermarked graphs. We generate ten BA graphs with density $\frac{|E|}{|V|}=5$ (according to the uniqueness experiment in Figure 2). For each of them, we first inserted a key $\omega$; then we regenerated five more graphs with the same generative parameters, and tried to extract $\omega$ from them. Extractions occur for different values of $\theta$. Figure 3 shows that as long as $\theta$ is less than or equal to 555, there are no false positives. Then, values of $\theta$ above 666 introduce some. Thus, $\theta=1.4$ is consistent with the 222-norm of the watermark to tolerate 101010 % edge flips for a graph of density 555 in BA models. Goal 3 is then achievable by parameterization.
In the remaining experiments, $\theta$ allows up to 101010 % edge flips since no false positives or false negatives occur in this setup.